OpenStudy (_malice_):

ΔRST is dilated about the origin, O, to create ΔR'S'T'. Point R is located at (3, 9), and point R' is located at (1.2, 3.6). Which scale factor was used to perform the dilation?

OpenStudy (zootedyoungin):

To find the scale factor, divide the first coordinates by the second coordinates: x-coordinates: 3÷1.2 =2.5 y-coordinates: 9÷3.6=2.5 The scale factor is: 2.5

OpenStudy (_malice_):

Incorrect. The scale factor is a ratio of the x- and y-coordinates in the image to the x- and y-coordinates in the preimage. Here, the scale factor is 0.4.

OpenStudy (zootedyoungin):

gg

OpenStudy (_malice_):

The scale factor is the ratio of R'O¯¯¯¯¯¯ to RO¯¯¯¯¯, where O is the origin.

OpenStudy (zootedyoungin):

oh

OpenStudy (zootedyoungin):

1.2/3 = 0.4 3.6/9 = 0.4

OpenStudy (_malice_):

:P

OpenStudy (_malice_):

medal anyways :P

OpenStudy (zootedyoungin):

yep

OpenStudy (_malice_):

thx peace